# Unsupervised Meta-Learning via In-Context Learning

### v1

@@ 4.2 Training details
We report the results following the $N$-way $K$-shot classification task typical of meta-learning algorithms, where $N=5$ and $K=1$ or $K=5$. All the models are trained for $100$ epochs, $500$ episodes per epoch, and $100$ fine-tuning steps at test time if required by the method. For CAMeLU, we choose not to apply any fine-tuning strategy to prove the strength of the training stage that does not require additional parameter updates during the inference phase. To validate the cross-domain performance, we trained the models on ImageNet-964 after removing the image labels. This is a variant of the original ImageNet-1k [8] where classes belonging to the validation and test splits of miniImageNet are removed. This is done to prevent any data leakage when evaluating the model on miniImageNet - a problem that is not taken into consideration in previous studies [12, 19]. For CAMeLU, we use a feature extractor pre-trained using CLIP [34] on a ViT-B/16 architecture and a class encoder consisting of a single learnable layer. In Appendix 7.4, we also report the results with different feature extractors. The transformer encoder consists of 8 encoder layers, each incorporating a multi-head self-attention block with 8 attention heads, an MLP with a reverse bottleneck of $3072$ and an input-output feature size of $1024$, where the last $768$ features are the ViT feature vector appended to the $256$ encoding of the labels. Finally, a single projection layer maps the transformer output to the predicted category. These choices were motivated by our hardware limitations, which constrained us to load and train the model on a consumer device with only 4GB of VRAM. The model is trained with the Adam optimizer with a learning rate of $10^{-5}$ and a warmup cosine scheduler as proposed in the original paper [46]. To account for statistical variations, each algorithm is run three times in full. The experiments are executed using Python and the PyTorch library on an Nvidia GeForce RTX 3050 Ti Laptop GPU with 4GB of VRAM. Further ablation experiments and competitors required a more capable architecture and are executed on an Nvidia A100-SXM4 GPU with 40GB of VRAM. More details about the training settings can be found in Appendix 7.1, while the code will be released upon acceptance of the paper.
@@ 4.3 Comparative results
Table 1 provides an overview of the experimental results for both the 5-way 1-shot and the 5-way 5-shot scenarios. The results demonstrate that CAMeLU outperforms the existing UML methods, regardless of the difference in the evaluation setting. As highlighted in Section 4.1, CACTUs, UMTRA, and Meta-GMVAE are evaluated only in-domain, requiring knowledge about the test domain prior to training. This is not necessary for CAMeLU as it demonstrates high performance in the challenging cross-domain scenario. Even compared to PsCo, the only UML method designed for cross-domain applications, CAMeLU exhibits a performance improvement in all datasets except miniImageNet, where PsCo achieves even better results than the supervised CAML. This is expected, as the model utilized for cross-domain evaluation of PsCo is pre-trained on ImageNet-1k, which includes the images of miniImageNet, leading to data leakage. Furthermore, PsCo requires a fine-tuning phase to adapt to the test domain, whereas CAMeLU achieves good performance with a single forward pass, enhancing its applicability to real-time applications. It is also worth noting that CAMeLU achieves comparable performance to its supervised counterpart, CAML, when evaluated on miniImageNet and CIFAR-fs and it even outperforms CAML when evaluated on more dissimilar domains, such as CUB and Aircraft. This highlights the efficacy of the task construction strategy used in CAMeLU, which acts as a sort of task augmentation and enhances the generalization capability of the model.
@@
Table 1: Performance comparison on miniImageNet, CIFAR-fs, CUB, and Aircraft dataset for 5-way 1-shot and 5-way 5-shot scenarios. In-domain approaches are trained with a Conv5 architecture, while PsCo cross-domain is trained on a ResNet-50 architecture. The symbol $\dagger$ indicates results that are affected by data leakage. The bold font highlights the best performing UML approach for each setting. Results show the mean and standard deviations across three complete runs of the algorithms.
@@
miniImageNet CIFAR-fs CUB Aircraft Method 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s In-Domain CACTUs-MAML $43.30\pm 0.29$ $54.21\pm 1.00$ $42.00\pm 1.47$ $56.64\pm 0.49$ $31.19\pm 0.37$ $36.81\pm 0.68$ $24.06\pm 0.78$ $27.26\pm 0.04$ CACTUs-ProtoNet $48.85\pm 0.69$ $62.52\pm 0.71$ $50.90\pm 0.46$ $64.52\pm 0.94$ $33.93\pm 0.37$ $44.41\pm 1.31$ $26.27\pm 0.28$ $30.88\pm 0.51$ UMTRA $39.93\pm 1.15$ $50.73\pm 0.67$ $32.93\pm 1.68$ $46.13\pm 2.81$ $27.06\pm 1.41$ $36.6\pm 2.43$ $22.40\pm 3.42$ $31.73\pm 2.25$ Meta-GMVAE $55.38\pm 0.90^{\dagger}$ $65.10\pm 0.64^{\dagger}$ $52.02\pm 0.88$ $64.18\pm 0.62$ $33.59\pm 0.63$ $39.09\pm 0.57$ $24.83\pm 0.51$ $27.60\pm 0.52$ PsCo $45.80\pm 0.01$ $62.64\pm 0.01$ $51.62\pm 0.01$ $68.43\pm 0.01$ $35.90\pm 0.01$ $47.46\pm 0.01$ $26.10\pm 0.01$ $31.40\pm 0.01$ Cross-Domain PsCo $\mathbf{82.08\pm 0.42}^{\dagger}$ $\mathbf{94.61\pm 0.15}^{\dagger}$ $\mathbf{66.28\pm 0.50}$ $83.36\pm 0.34$ $52.16\pm 0.53$ $75.35\pm 0.44$ $27.17\pm 0.32$ $42.01\pm 0.43$ CAMeLU $69.63\pm 2.87$ $90.95\pm 0.54$ $58.60\pm 1.92$ $\mathbf{84.26\pm 0.77}$ $\mathbf{55.12\pm 1.42}$ $\mathbf{77.71\pm 3.17}$ $\mathbf{36.93\pm 0.53}$ $\mathbf{46.27\pm 2.18}$ CAML (supervised) $92.58\pm 0.12$ $97.93\pm 0.05$ $66.33\pm 3.22$ $83.39\pm 1.28$ $45.76\pm 0.81$ $68.40\pm 2.55$ $33.04\pm 1.82$ $35.82\pm 0.92$
@@ 4.4 Memorization to generalization phase shift
During the training of CAMeLU, we observed a distinct trend in the validation accuracy, similar to the findings in [22]. Figure 2 illustrates this pattern, showing the validation accuracy relative to its initial value, or, in other words, how much the model learns from datasets different from the one we are training on. Specifically, the curves in Figure 2 resemble a logistic curve, which can be divided into three phases that we denote as memorization, learning, and generalization. In the memorization phase, the model memorizes the tasks seen during training and extends this knowledge to unseen tasks, resulting in a slight improvement for datasets with high similarity with ImageNet-964 (e.g., miniImageNet). For the other datasets, instead, transferring this knowledge can even result in a performance decrease due to the intrinsic domain distance of the dataset (see CUB and Aircraft, which are fine-grained datasets). As training progresses and the model observes more tasks, the learning phase occurs. This phase is characterized by a transition to the learning-to-learn state where the model learns to identify the tasks and to extract the features that are more useful for solving them.
@@
The duration of this phase varies, with datasets like miniImageNet and CIFAR-fs exhibiting rapid learning within approximately $10$ epochs, while datasets such as CUB and Aircraft may necessitate up to 40 epochs. This timespan depends on several factors, including the similarity between the training and evaluation datasets, the size of the test dataset, and the model’s learning ability [22, 33]. For instance, CUB, with its fine-grained nature and small test set size (around $1770$ images) necessitates a longer learning phase compared to the miniImageNet dataset (which has a test set with 12 0001200012\,00012 000 images). Subsequently, in the generalization phase, the model can generalize to tasks significantly different from those observed during training using a single forward pass. Further analyses about the generalization capabilities of CAMeLU and the number of epochs required for reaching the generalization phase are presented in Appendices 7.5 and 7.6.
4.5 Multi-dataset training
Table 2: Comparison of CAMeLU and CAML when trained on the single ImageNet-964 dataset (denoted as IN-964) and on a multi-dataset consisting of ImageNet-964 + MSCOCO + Fungi. Results show the mean and standard deviations for the 5-way 1-shot and the 5-way 5-shot settings across three complete runs of the algorithms.
@@ 4.5 Multi-dataset training
miniImageNet CIFAR-fs CUB Aircraft 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s CAMeLU (IN-964) $69.63\pm 2.87$ $90.95\pm 0.54$ $58.60\pm 1.92$ $84.26\pm 0.77$ $55.12\pm 1.42$ $77.71\pm 3.17$ $36.93\pm 0.53$ $46.27\pm 2.18$ CAMeLU (mds) $70.42\pm 2.77$ $91.02\pm 0.97$ $60.86\pm 3.94$ $\mathbf{85.05\pm 0.27}$ $\mathbf{55.58\pm 2.68}$ $\mathbf{78.68\pm 2.13}$ $\mathbf{38.88\pm 2.22}$ $\mathbf{48.86\pm 3.45}$ CAML (IN-964) $\mathbf{92.58\pm 0.12}$ $\mathbf{97.93\pm 0.05}$ $66.33\pm 3.22$ $83.39\pm 1.28$ $45.76\pm 0.81$ $68.40\pm 2.55$ $33.04\pm 1.82$ $35.82\pm 0.92$ CAML (mds) $92.10\pm 0.33$ $97.82\pm 0.03$ $\mathbf{67.01\pm 2.66}$ $\mathbf{85.05\pm 2.88}$ $44.03\pm 2.93$ $73.47\pm 2.56$ $31.51\pm 0.58$ $38.88\pm 4.54$
We conduct additional experiments to evaluate how CAMeLU performs when trained on a large-scale dataset. As our focus is cross-domain classification through in-context learning, we hypothesize that training on a dataset spanning various concepts enhances classification performance, as suggested in [12, 31]. To this effect, we combine three training datasets with different levels of granularity: ImageNet-964 [8], MSCOCO [27], and Fungi [38]. At each training episode, a dataset is uniformly sampled, and $N$ data points are subsequently extracted. These samples are then utilized in our method as described in Section 3. Table 2 presents the results for CAMeLU and the supervised CAML method. Notably, training on a combination of multiple datasets (denoted as mds in Table 2) yields improved performance compared to training solely on ImageNet-964. However, the performance gap is not substantial, likely due to the large size of ImageNet-964 (around $80\%$ of the overall dataset), resulting in its more frequent selection compared to the other datasets. Therefore, to optimize computational resources and time, we conducted all the remaining experiments by training only on ImageNet-964.
@@ 4.6 Comparison with SSL methods
In this section, we compare CAMeLU with SimCLR [7] and SwAV [6]. For training SSL methods in our experiments, we employed a backbone network with a ResNet-50 architecture pre-trained on ImageNet-1k and obtained from PyTorch Lightning Bolts [3]. To facilitate model adaptation to the test domain, we fine-tuned a classification layer on top of the pre-trained backbone using SGD with initial learning rate of $0.01$, momentum of $0.9$, weight decay of $10^{-4}$, and $100$ fine-tuning steps per task, following [19]. This setup differs from the evaluation setting used in CAMeLU, where predictions are obtained with a single forward pass, leveraging the in-context learning ability of transformer architectures. However, SSL approaches must adapt to the test domain before label predictions, resulting in a less challenging evaluation setting than CAMeLU. Results for SSL approaches are averaged over $500$ test tasks and presented in Table 3. While SSL approaches outperform CAMeLU on miniImageNet and CIFAR-fs, their performance decreases when evaluated on the other datasets. CUB and Aircraft are two fine-grained datasets significantly different from ImageNet-1k, challenging the transferability of features learned by SSL methods to these datasets. Moreover, the high performance on miniImageNet may be attributed to data leakage with ImageNet-1k, as discussed in Section 4.2. CAMeLU, in contrast, demonstrates effective generalization to tasks sampled from these datasets, once again highlighting its generalization ability over mere memorization.
@@
Table 3: Comparison between CAMeLU and SSL approaches for the 5-way 1-shot and 5-way 5-shot scenario on miniImageNet, CIAFR-fs, CUB, and Aircraft. The symbol $\dagger$ indicates results that are affected by data leakage. Results show the mean and standard deviations across three complete runs of the algorithms.
@@
miniImageNet CIFAR-fs CUB Aircraft Method 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s SimCLR $\mathbf{83.32\pm 0.23}^{\dagger}$ $\mathbf{94.86\pm 0.61}^{\dagger}$ $\mathbf{64.52\pm 0.69}$ $84.36\pm 0.40$ $47.35\pm 0.53$ $66.87\pm 0.82$ $29.36\pm 0.90$ $39.99\pm 0.86$ SwAV $74.83\pm 0.71^{\dagger}$ $\mathbf{94.96\pm 0.91}^{\dagger}$ $66.97\pm 0.15$ $\mathbf{87.14\pm 0.10}$ $47.84\pm 0.31$ $69.31\pm 0.01$ $30.33\pm 0.31$ $\mathbf{47.43\pm 0.11}$ CAMeLU $69.63\pm 2.87$ $90.95\pm 0.54$ $58.60\pm 1.92$ $84.26\pm 0.77$ $\mathbf{55.12\pm 1.42}$ $\mathbf{77.71\pm 3.17}$ $\mathbf{36.93\pm 0.53}$ $46.27\pm 2.18$
5 Limitations
Despite the promising results demonstrated by our novel UML approach on several datasets, some limitations remain. Its applicability and robustness in real-world scenarios with diverse and noisy data remain to be thoroughly evaluated. In real-world applications, data can be incomplete, mislabeled, or drawn from significantly different distributions, leading to potential degradation in the model’s performance in the presence of noisy or corrupted data. Additionally, CAMeLU’s performance depends on the pre-trained feature extractor, as demonstrated in Appendix 7.4. Replacing the extractor pre-trained with CLIP with one pre-trained using an SSL strategy could make the pipeline completely unsupervised, albeit at the price of performance degradation. Lastly, the proposed approach is designed to handle a fixed number of classes (ways) per task during training and testing, requiring knowing the value of $N$ in advance. Modifications to the task creation and training process would be necessary to extend our approach to handle an arbitrary number of ways.

### v2

@@ 4.2 Training details
We report the results following the $N$-way $K$-shot classification task typical of meta-learning algorithms, where $N=5$ and $K=1$ or $K=5$. All models are trained for $100$ epochs, $500$ episodes per epoch, and $100$ fine-tuning steps at test time if required by the method. For CAMeLU, we choose not to apply any fine-tuning strategy to prove the strength of the training stage that does not require additional parameter updates during the inference phase. Furthermore, we introduce ImageNet-964, a variant of the original ImageNet-1k dataset [8] where the classes belonging to the validation and test splits of miniImageNet are removed to prevent data leakage when evaluating the model on miniImageNet - a problem that is not taken into consideration by previous studies [12, 20]. To provide a fair comparison, all cross-domain methods are trained using ImageNet-964. For CAMeLU, we use a feature extractor pre-trained using ResNet-50 [18] on ImageNet-964 and a class encoder consisting of a single learnable layer. In Appendix A.4, we also report the results with different feature extractors. The transformer encoder consists of 8 encoder layers, each incorporating a multi-head self-attention block with 8 attention heads, an MLP with a reverse bottleneck of $3072$ and an input-output feature size of $2304$, where the last $2048$ features is the ResNet-50 feature vector appended to the $256$ encoding of the labels. Finally, a single projection layer maps the transformer output to the predicted category. The model is trained with the Adam optimizer with a learning rate of $10^{-5}$ and a warmup cosine scheduler as proposed in the original paper [48]. To account for statistical variations, each algorithm is run three times in full and the complete results reporting the standard deviations are presented in Appendix A.10. The experiments are executed using Python and the PyTorch library on an Nvidia GeForce RTX 3070 Ti Laptop GPU with 8GB of VRAM. Further ablation experiments and competitors required a more capable architecture and are executed on an Nvidia A100-SXM4 GPU with 40GB of VRAM. More details about the training settings can be found in Appendix A.1, while the code will be released upon acceptance of the paper.
@@ 4.3 Comparative results
Table 1: Performance comparison on miniImageNet, CIFAR-fs, CUB, Aircraft, and Meta-iNat datasets for 5-way 1-shot and 5-way 5-shot scenarios. Cross-domain approaches are trained using ImageNet-964 and a ResNet-50 feature extractor. The symbol $\dagger$ indicates results that are affected by data leakage. The bold font highlights the best performing UML approach for each setting. Results show the average across three complete runs of the algorithms. Complete results with standard deviations are reported in Table 10 in Appendix A.10.
@@
miniImageNet CIFAR-fs CUB Aircraft Meta-iNat Method 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s In-Domain CACTUs-MAML $43.30$ $54.21$ $42.00$ $56.64$ $31.19$ $36.81$ $24.06$ $27.26$ $20.13$ $21.84$ CACTUs-ProtoNet $48.85$ $62.52$ $50.90$ $64.52$ $33.93$ $44.41$ $26.27$ $30.88$ $27.30$ $29.08$ UMTRA $39.93$ $50.73$ $32.93$ $46.13$ $27.06$ $36.6$ $22.40$ $31.73$ $28.96$ $37.12$ Meta-GMVAE $55.38^{\dagger}$ $65.10^{\dagger}$ $52.02$ $64.18$ $33.59$ $39.09$ $24.83$ $27.60$ $34.22$ $40.23$ PsCo $47.29$ $64.85$ $42.21$ $62.92$ $33.09$ $51.02$ $26.19$ $38.80$ $36.97$ $55.88$ Cross-Domain PsCo $67.89$ $90.17$ $53.34$ $76.22$ $43.35$ $70.19$ $29.87$ $38.20$ $46.21$ $70.05$ CAMeLU $\mathbf{76.51}$ $\mathbf{92.14}$ $\mathbf{61.79}$ $\mathbf{80.43}$ $\mathbf{65.52}$ $\mathbf{80.35}$ $\mathbf{33.17}$ $\mathbf{39.11}$ $\mathbf{57.27}$ $\mathbf{75.45}$ CAML (supervised) $81.75$ $92.31$ $59.44$ $75.27$ $54.63$ $66.81$ $28.92$ $32.06$ $50.86$ $67.07$
Table 1 provides an overview of the experimental results for both the 5-way 1-shot and the 5-way 5-shot scenarios. The results demonstrate that CAMeLU outperforms the existing UML methods, regardless of the difference in the evaluation setting. As highlighted in Section 4.1, CACTUs, UMTRA, and Meta-GMVAE are evaluated only in-domain, requiring knowledge about the test domain prior to training. This is not necessary for CAMeLU as it demonstrates high performance in the challenging cross-domain scenario. Even compared to PsCo, the only UML method designed for cross-domain applications, CAMeLU exhibits a performance improvement in all datasets. Furthermore, PsCo requires a fine-tuning phase to adapt to the test domain, whereas CAMeLU achieves good performance with a single forward pass, enhancing its applicability to real-time applications. It is also worth noting that CAMeLU achieves comparable performance to its supervised counterpart, CAML, when evaluated on miniImageNet and it even outperforms CAML when evaluated on more dissimilar domains, such as CUB, Aircraft, and Meta-iNat. This finding highlights the efficacy of the task construction strategy used in CAMeLU, which acts as a sort of task augmentation and enhances the generalization capability of the model.
@@ 4.4 Memorization to generalization phase shift
During the training of CAMeLU, we observed a distinct trend in the validation accuracy, similar to the findings in [23]. Figure 2 illustrates this pattern, showing the validation accuracy relative to its initial value, or, in other words, how much the model learns from datasets different from the one we are training on. Specifically, the curves in Figure 2 resemble a logistic curve,
@@
which can be divided into three phases that we denote as memorization, learning, and generalization. In the memorization phase, the model memorizes the tasks seen during training and extends this knowledge to unseen tasks, resulting in a slight improvement for datasets with high similarity with ImageNet-964 (e.g., miniImageNet and CIFAR-fs). For the other datasets, instead, transferring this knowledge can even result in a performance decrease due to the intrinsic domain distance of the dataset (see CUB and Aircraft, which are fine-grained datasets). As training progresses and the model observes more tasks, the learning phase occurs. This phase is characterized by a transition to the learning-to-learn state where the model learns to identify the tasks and to extract the features that are more useful for solving them. The duration of this phase varies, with datasets like miniImageNet and CIFAR-fs exhibiting rapid learning within approximately $10$ epochs, while datasets such as CUB and Aircraft may necessitate up to 40 epochs. This timespan depends on several factors, including the similarity between the training and evaluation datasets, the size of the test dataset, and the model’s learning ability [23, 35]. For instance, CUB, with its fine-grained nature and small test set size (around $1770$ images), necessitates a longer learning phase compared to the miniImageNet dataset (which has a test set with 12 0001200012\,00012 000 images). Subsequently, in the generalization phase, the model can generalize to tasks significantly different from those observed during training using a single forward pass. Further analyses about the generalization capabilities of CAMeLU and the number of epochs required for reaching the generalization phase are presented in Section 4.5 and Appendix A.8.
4.5 Generalization on small-scale datasets
Table 2: Accuracy results obtained training PsCo, BECLR, and CAMeLU with a small-scale dataset, namely miniImageNet, denoted as (mini) in the table. Results show both in-domain performance (on the test set of miniImageNet) and cross-domain performance on CIFAR-fs, CUB, Aircraft, and Meta-iNat. The average results across three complete runs of the algorithms are reported. Complete results with standard deviations are presented in Table 11 in Appendix A.10.
@@ 4.5 Generalization on small-scale datasets
miniImageNet CIFAR-fs CUB Aircraft Meta-iNat 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s PsCo (mini) $47.29$ $64.85$ $42.21$ $62.92$ $33.09$ $51.02$ $26.19$ $\mathbf{38.80}$ $36.97$ $55.88$ BECLR (mini) $\mathbf{81.04}$ $87.88$ $57.05$ $72.82$ $42.47$ $58.03$ $27.48$ $38.46$ $49.87$ $65.05$ CAMeLU (mini) $75.99$ $\mathbf{90.38}$ $\mathbf{61.25}$ $\mathbf{78.79}$ $\mathbf{60.60}$ $\mathbf{74.77}$ $\mathbf{31.39}$ $36.52$ $\mathbf{55.60}$ $\mathbf{72.12}$
While most studies on training transformer architectures focus on large-scale training datasets, we investigate the generalization capabilities of CAMeLU using a small-scale training dataset. Specifically, we train CAMeLU on miniImageNet and evaluate its performance both in-domain (i.e., on the test set of miniImageNet) and cross-domain on CIFAR-fs, CUB, Aircraft, and Meta-iNat. CAMeLU demonstrates effective generalization in this scenario, showing impressive performance in both in-domain and cross-domain settings, as shown in Table 2, surpassing PsCo and BECLR by a large margin.
Additionally, a comparison of CAMeLU’s performance when trained on a small-scale dataset like miniImageNet (Table 2), on ImageNet-964 (Table 1), and on a large-scale dataset (Table 5 in Appendix A.3) show that our method is only slightly affected by the size of the training dataset.
Figure 3: Relative validation accuracy comparison between CAMeLU (orange) and CAML (blue) when evaluated in-domain on miniImagenet. The relative validation accuracy is computed as described in Section 4.4. The curve obtained with CAMeLU reflects the three phases of memorization, learning, and generalization even when using a small-scale dataset.
This robustness enhances CAMeLU’s applicability to scenarios where only a small unlabeled training dataset is available, which is common in real-world applications.
Finally, Figure 3 shows the relative validation accuracy of CAMeLU and CAML while trained and evaluated on miniImageNet. While the curve obtained with CAMeLU reflects the three phases of memorization, learning, and generalization discussed in Section 4.4, the relative validation accuracy of CAML remains flat. This difference may be attributed to the task creation mechanism of CAMeLU, which acts as a task augmentation strategy, increasing the variability of tasks presented to the model during training and thereby enhancing its generalization capabilities.
@@ 4.6 Comparison with SSL methods
Table 3: Comparison between CAMeLU and SSL approaches for the 5-way 1-shot and 5-way 5-shot scenario on miniImageNet, CIFAR-fs, CUB, Aircraft, and Meta-iNat. The symbol $\dagger$ indicates results that are affected by data leakage. Results show the average across three complete runs of the algorithms. Complete results with standard deviations are reported in Table 12 in Appendix A.10.
@@
miniImageNet CIFAR-fs CUB Aircraft Meta-iNat Method 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s SimCLR $\mathbf{83.32}^{\dagger}$ $94.86^{\dagger}$ $64.52$ $84.36$ $47.35$ $66.87$ $29.36$ $39.99$ $52.44$ $73.19$ SwAV $74.83^{\dagger}$ $\mathbf{94.96}^{\dagger}$ $\mathbf{66.97}$ $\mathbf{87.14}$ $47.84$ $69.31$ $30.33$ $\mathbf{47.43}$ $53.57$ $74.53$ CAMeLU $76.51$ $92.14$ $61.79$ $80.43$ $\mathbf{65.52}$ $\mathbf{80.35}$ $\mathbf{33.17}$ $39.11$ $\mathbf{57.27}$ $\mathbf{75.45}$
In this section, we compare CAMeLU with SimCLR [7] and SwAV [6]. For training SSL methods in our experiments, we employed a backbone network with a ResNet-50 architecture pre-trained on ImageNet-1k obtained from PyTorch Lightning Bolts [3]. While this setup leads to data leakage when evaluated on miniImageNet, due to overlap between the test and training classes, pre-training these SSL approaches from scratch using a different training dataset was beyond our available computational resources. To facilitate model adaptation to the test domain, we fine-tuned a classification layer on top of the pre-trained backbone using SGD with initial learning rate of $0.01$, momentum of $0.9$, weight decay of $10^{-4}$, and $100$ fine-tuning steps per task, following [20]. This setup differs from the evaluation setting used in CAMeLU, where predictions are obtained with a single forward pass, leveraging the in-context learning ability of transformer architectures. However, SSL approaches must adapt to the test domain before label predictions, resulting in a less challenging evaluation setting than CAMeLU. Results for SSL approaches are averaged over $500$ test tasks and presented in Table 3. While SSL approaches outperform CAMeLU on miniImageNet and CIFAR-fs, their performance decreases when evaluated on the other datasets. CUB, Aircraft, and Meta-iNat are fine-grained datasets significantly different from ImageNet-1k, challenging the transferability of features learned by SSL methods to these datasets. Moreover, the high performance on miniImageNet and CIFAR-fs may be attributed to the presence of data leakage with ImageNet-1k and the high similarity with CIFAR-fs, as discussed in Section 4.2. CAMeLU, in contrast, demonstrates effective generalization to tasks sampled from these datasets, once again highlighting its generalization ability over mere memorization.